OpenStudy (anonymous):

PLEASE HELP ME (kind of in a hurry) Choose the correct answer. Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? Joelle: 270, Anita: 155 Joelle: 180, Anita: 245 Joelle: 140, Anita: 240 Joelle: 155, Anita: 270

OpenStudy (anonymous):

155 270

OpenStudy (anonymous):

thanks !

OpenStudy (kamibug):

a + j = 425 a = 2j - 40 2j - 40 + j = 425 3j - 40 = 425 3j = 465 j = 155 <--- a + 155 = 425 a = 270 <---

OpenStudy (anonymous):

155 to 270